Find the unknown quantity in the given equation.
log3 2x - log3 (x + 5) = 0

Answer by Alan3354(69443)

Find the unknown quantity in the given equation.
log3 2x - log3 (x + 5) = 0
----
log3 2x = log3 (x + 5)
2x = x+5
etc